OpenStudy (anonymous):

What is an equation of the line that passes through the point (3, 6) and has a slope of -3/4?

myininaya (myininaya):

do you know point slope formula?

OpenStudy (anonymous):

yes its y-y1=m(x-x1) right?

myininaya (myininaya):

so what will you replace m with?

OpenStudy (anonymous):

-3/4

myininaya (myininaya):

great what about x1?

OpenStudy (anonymous):

3

myininaya (myininaya):

and lastly y1?

OpenStudy (anonymous):

6

myininaya (myininaya):

great

myininaya (myininaya):

when you make all of those replacements in that equation you gave above you have found the equation of the line in point slope form

OpenStudy (anonymous):

i need it in slope intercept form so how do I convert it?

myininaya (myininaya):

well if you have something writen like this y-y1=m(x-x1) and you want y=mx+b then you will need to get the constant y1 on the other side so adding y_1 on both sides gives y=m(x-x1)+y1 now distribute the m like so y=mx-mx1+y1 y=mx+(-mx1+y1) where m is the slope and the y-intercept is -mx1+y1

myininaya (myininaya):

if you don't like that because you have trouble with there not being actual number there you can write the equation you already have and i can help you get to that form

OpenStudy (anonymous):

for the point slope form i got y= 5 and 1/4x + 8 and 1/4

myininaya (myininaya):

what do the and's symbolize you should have something in this form y-y1=m(x-x1)

myininaya (myininaya):

you gave me the numbers to put in

myininaya (myininaya):

you said m was -3/4 x1 was 3 y1 was 6

myininaya (myininaya):

\[y-6=\frac{-3}{4}(x-3)\]

myininaya (myininaya):

in my general example follow the steps I took to put into y=mx+b form add the y1 on both sides (in this case add the 6 on both sides)

myininaya (myininaya):

next i distribute the m to the terms in the parenthesis so in this case you multiply -3/4 to both terms in the parenthesis

myininaya (myininaya):

can you do this and show me what you get

OpenStudy (anonymous):

i get -3/4x+6/3/4

myininaya (myininaya):

so you did this \[y=\frac{-3}{4}(x-3)+6\] \[y=\frac{-3}{4}x+\frac{3}{4}(3)+6\]

myininaya (myininaya):

the only thing left to do is do 3/4(3)+6

myininaya (myininaya):

what is 3 times 3/4?

OpenStudy (anonymous):

2/1/4

myininaya (myininaya):

ok or 9/4

myininaya (myininaya):

then what is 9/4+6?

OpenStudy (anonymous):

8/1/4

myininaya (myininaya):

ok or 33/4

myininaya (myininaya):

so remember we just simplify the y-intercept part which was 3*3/4+6

OpenStudy (anonymous):

so all together it would be y=-3/4+33/4?

myininaya (myininaya):

don't forget the x that is suppose to be next to the -3/4

myininaya (myininaya):

awww... thanks rob

OpenStudy (anonymous):

A medal for @myininaya's patience.

OpenStudy (anonymous):

ok thank you so much you were a big help!!

myininaya (myininaya):

thanks @Its_ian2000

myininaya (myininaya):

i try